[Brahms-l] Shift report 20040402 00:00-08:00

From: Apache <apache@bnl.gov>
Date: Fri Apr 02 2004 - 07:05:01 EST
------------------------------------------------------------
Date: Fri Apr  2 07:05:00 2004, Shift: 00:00-08:00
Supervisor: sl, with: 
Log book#15 pages 105 - 105

Shift Summary:
------------------------------------------------------------
1.One good last store that was kept for 6 hours.
2.Data run:11565-11568, 11570, 11575.
~100k tr3 events at 90 deg 6kG A.
~230k tr6 events at 8 deg A pol.(Only FFS)
3.D3 still out of function.
  

------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Fri Apr 2 07:05:13 2004

This archive was generated by hypermail 2.1.8 : Fri Apr 02 2004 - 07:05:30 EST